# Towards General Industrial Intelligence: A Survey on Industrial IoT-Enhanced Continual Large Models

### v1

@@ IV-A3 Parameter Isolation Methods
Parameter Isolation Methods learn specific parameters within a network for each task to prevent catastrophic forgetting. Depending on whether the network structure remains fixed during training, parameter isolation methods can be categorized into static and dynamic architecture.
Static Architecture Methods. Static architecture methods allocate their fixed parameters in limited network architecture to tasks for task-specific learning. PathNet [175] introduces pathways containing active modules in each layer through a network. Each task is assigned its optimal pathway. Once task $t$ has been trained, its optimal fit pathway becomes fixed, preventing further parameter changes. PackNet [176] utilizes pruning to free up parameters for new tasks. During the training of PackNet, a network is learned for the initial task. Next, a portion of the learned weights is set to zero for pruning, and these pruned weights are then learned for the second task. In subsequent training, after the network has learned task $t$, pruning is applied to the weights acquired for that task. The pruned weights then undergo retraining for task $t+1$, while the weights related to tasks prior to task $t+1$ remain frozen to retain previously acquired knowledge. Masks are stored to indicate the relevant active parameters when the network learns a new task. HAT [177] introduces a task-based hard attention mechanism with unit-based masks to retain the previous knowledge and constrain updates for news tasks. More recently, DSN [182] proposes a neuron-wise mask mechanism to select the affiliated weights of a small set of neurons for previous knowledge transfer.
Dynamic Architecture Methods. Dynamic architecture methods expand the network as new tasks are learned while retaining the parameters for previous tasks. PNNs [183] adapt to each task by creating a new network for each one. Lateral connections to features of previous networks are used to transfer previous knowledge. DEN [225] utilizes group-sparsity regularization to determine the addition of neurons at each layer for each task, thereby avoiding unnecessary network retraining. RCL [226] trains a recurrent neural network to determine the optimal architectural hyper-parameters for each task. The recurrent neural network is guided by a reward signal derived from reinforcement learning. Based on Mixture-of-Expert [227], some methods [184, 191] utilize task-specific experts for network expansion. In Expert Gate [184], a new expert model is integrated each time a new task needs to be learned, with knowledge transferred from prior models. Dynamic architecture approaches are also employed in continual graph learning [186, 187].
IV-B Pre-trained Models Based CL
In this section, we will explore Pre-trained Models (PTMs) based CL. Inspired by [228], we further summarize and introduce developments in PTM-based continual learning.
Figure 26: Mixture of Experts.
IV-B1 Mixture of Experts-based Methods
As shown in Fig. 26, the Mixture of Experts (MoE) employs a modular approach to enhance both performance and efficiency in neural networks. This approach consists of multiple expert sub-networks, each specializing in different tasks, coordinated by a routing network. The routing network directs incoming data to the most relevant experts based on data characteristics. These experts process the data independently, and their outputs are often combined through mechanisms like weighted averaging to generate the final response.
The modular structure of MoE reduces the overall parameter load by activating only the necessary experts for each input, making it well-suited for large-scale, diverse datasets. In CL settings, MoE helps mitigate catastrophic forgetting by dynamically adjusting which experts are involved based on the task at hand [229].
Recent CL advancements have integrated MoE with foundational models to boost adaptability and minimize forgetting. For example, MoE adapters in pre-trained models such as CLIP [20] enable incremental learning without the high computational costs usually associated with extensive model updates. These adapters dynamically adjust to new tasks, ensuring sustained performance across evolving learning scenarios.
Innovations like the expert aggregation loss mechanism further enhance model adaptability by fine-tuning expert contributions in real-time, based on confidence levels and specific knowledge requirements, providing an efficient framework for managing diverse data streams [230, 231].
IV-B2 Analytic Learning-based Methods
Figure 27: Analytic CL Based on Pre-trained Models vs. Traditional Replay-based CL. Analytic CL, which is example-free, demonstrates performance comparable to that of replay-based methods.
Training deep models using backpropagation has inherent limitations, such as gradient vanishing/exploding, divergence during iterations, and long training cycles. Analytic Learning (AL) addresses these issues by employing recursive least squares (see Fig.  27). Utilizing the matrix inversion principle, its effectiveness has been preliminarily validated. However, since the weights are learned by processing the entire dataset, AL may encounter memory insufficiency problems. This issue can be mitigated by Block-wise Recursive Moore-Penrose Learning (BRMP) [232], which replaces joint learning with a recursive approach while maintaining equivalent model accuracy.
Analytic Class Incremental Learning (ACIL) [233] is the first method to apply Analytic Learning to Class Incremental Learning (CIL) problems. This method is example-free, providing an effective strategy to counter catastrophic forgetting and reducing resource consumption while achieving performance comparable to replay-based methods. Gaussian Kernel Embedded Analytic Learning [234] further addresses catastrophic forgetting in few-shot scenarios by employing kernel methods.
IV-B3 Prompt-based Methods
VPT [111] applies a prompt-based tuning approach where a set of trainable prompts $P=\{P_{1},P_{2},\ldots,P_{M}\}$ is introduced at the model’s input frontend. Combined with input data, these prompts directly influence the forward propagation process to encode task-relevant information.
The selection and application of prompts are realized through an optimization problem, formulated as:
$\min_{P,W}\sum_{(x,y)\in D}\ell(W^{T}\phi(x;P),y),$ (19)
where $\phi(x;P)$ represents the prompted input features, and $\ell$ is the loss function measuring the deviation between predictions and actual values.
Prompt-based CL methods provide an efficient way to adjust pre-trained models, enabling them to adapt quickly to new tasks while retaining previously learned knowledge. A prompt pool is designed to manage and apply these prompts [114], allowing the model to choose the most suitable prompts for specific instances. The selection process is governed by a retrieval function $g(x)$, which evaluates each prompt’s relevance to the current input $x$. The selection strategy can be expressed as:
$K_{x}=\operatorname{argmin}_{s_{i}\in[1,M],i=1}^{N}\gamma(\phi(x,k_{s_{i}})),$ (20)
where $k_{s_{i}}$ are keys associated with the prompts $P_{s_{i}}$, and $\gamma$ represents the cosine distance, used to assess the similarity between input features and keys.
IV-B4 Representation-based Methods
Pre-trained models are widely used for classifying new tasks due to their strong representational abilities. However, evaluating their inherent classification capacity poses challenges. SimpleCIL [235] offers a straightforward and effective approach by calculating the center or prototype of each class in the data stream:
$c_{i}=\frac{1}{K}\sum_{j=1}^{|D^{b}|}\mathbb{I}(y_{j}=i)\phi(x_{j}),$ (21)
where $\mathbb{I}(\cdot)$ is the indicator function which outputs 1 if the expression holds and 0 otherwise, $K$ is the number of samples for class $i$, and $\phi(x_{j})$ is the embedding of input $x_{j}$. This method extracts each class’s most common representational pattern, providing a strong starting point for subsequent classification tasks.
ADAM [236] enhances adaptability by linking the features of pre-trained models with those of task-specific fine-tuned models. This approach bridges domain gaps, improving the generalizability and task adaptability of representations through parameter-efficient modules:
$c_{i}=\frac{1}{K}\sum_{j=1}^{|D^{b}|}\mathbb{I}(y_{j}=i)[\phi(x_{j}),\phi(x_{j};\text{PEFT})],$ (22)
where $\phi(x_{j};\text{PEFT})$ represents the task-specific fine-tuned model embeddings.
To further enhance classification performance, RanPAC [237] introduces feature projection into a high-dimensional space to increase class separability:
$c_{i}=\frac{1}{K}\sum_{j=1}^{|D^{b}|}\mathbb{I}(y_{j}=i)\text{Proj}(\phi(x_{j};\text{PEFT})).$ (23)
The methods discussed in this section demonstrate how to effectively utilize the inherent representational capabilities of pre-trained models to adapt to new tasks and environments, which is crucial for deploying CL systems in dynamic settings.
IV-C Efficient Inference Algorithm
Figure 28: Test Time Adaptation.
IV-C1 Test Time Adaptation
As shown in Fig. 28, Test-time Adaptation (TTA) adjusts a pre-trained model to unlabeled data from a target domain at the time of testing, differentiating it from Domain Generalization which only operates during training. TTA enhances model performance by utilizing available test data, unlike Domain Adaptation that requires access to both source and target domain data. This makes TTA especially suitable for contexts where data privacy is paramount. TTA methodologies can be divided into test-time domain adaptation, test-time batch adaptation, and online test-time adaptation, based on test data characteristics.
The Edge-Cloud Learning Model [238] adapts to dynamic edge environments through a modular architecture. This approach enables the derivation of personalized edge models from large cloud-based models, tailored to specific needs while minimizing parameter conflicts. This modular setup allows flexible and efficient model adaptation in resource-constrained environments.
Continual Test-Time Adaptation addresses the challenge of adapting models to continuously evolving target domains. The introduction of the Visual Domain Adapter (ViDA) [239] allows the model to handle both domain-specific and shared knowledge. The Homeostatic Knowledge Allotment strategy optimizes the integration and utilization of knowledge from different ViDA, enhancing the model’s adaptability and preserving its plasticity across varying domains.
IV-C2 Edge Cloud Collaboration
Resource constraints in edge environments often hinder local deployment of LMs, leading to significant inference delays [240] that challenge real-time IIoT intelligent perception applications such as autonomous driving, metaverse [241], and digital twins [10]. Edge-cloud collaborative IIoT perception methods, leveraging cooperative inference between large and small models, provide a viable solution to this issue [242, 243]. By employing a hard example mining algorithm, difficult samples are selectively sent to the cloud for processing by a large visual model, while simpler samples are handled locally by small models, thereby enhancing performance on challenging samples while ensuring low inference latency.
As illustrated in Fig. 29, in edge-cloud collaborative intelligent driving systems, cloud-based large models play a critical role, particularly in navigating complex environments with high object density or when local model performance is compromised by visual impairments or lighting variations. Leveraging cloud resources for deeper analysis enhances accuracy, reliability, and safety in these systems.
The Cloud-Edge Collaborative Computing (CECC) framework [244] integrates computing resources from network back-end and edge nodes to deliver more flexible services, striking an optimal balance between resource abundance and responsiveness. Strategies aimed at mitigating edge device limitations emphasize enhancing edge-cloud collaboration. For example, the EdgeFM system [245] augments edge capabilities using cloud servers for querying cloud-based Feature Maps, ensuring real-time model adaptation and maintaining high accuracy under dynamic operational conditions. The EVOLVE framework [230] further advances cloud-edge synergy by employing multiple cloud-deployed expert models to refine self-supervised learning on edge devices. Featuring a dynamic expert aggregation loss mechanism, EVOLVE adjusts model weights in real-time based on reliability and expertise, effectively managing dynamic data streams and sustaining high accuracy in dynamic environments.
Figure 29: Edge-Cloud Collaborative Driving Planning with LMs.
IV-C3 Federated Learning
Research further suggests that the integration of FL with GII holds considerable promise for future advancements [250]. The combination of FL with the IIoT expands application possibilities across multiple domains, such as data sharing, task offloading, attack detection, localization, and privacy/security enhancement [33, 251].
In particular, FL-based approaches have demonstrated significant potential in critical sectors like smart healthcare, smart transportation, unmanned aerial vehicles (UAVs), smart cities, and intelligent manufacturing. These areas benefit from the enhanced data privacy and distributed computing capabilities enabled by FL, highlighting its role in supporting the complex infrastructures required for next-generation IIoT systems. The practical applications of FL in these fields demonstrate its potential to facilitate more secure, efficient, and scalable solutions for IIoT-driven industrial systems.
IV-D Benchmarks
Benchmarking is crucial for evaluating and improving LLMs. “Continual Instruction Tuning” [231] was proposed as a new evaluation benchmark to assess how machine learning models maintain existing skills while incorporating new knowledge. This benchmark uses ten diverse datasets covering eight task types to test models on their instruction adherence and general knowledge retention, effectively measuring their alignment with human intentions and reasoning abilities.
The SPHINX-X series was developed to enhance the SPHINX framework for increased architectural and training efficiency [252]. This series streamlines the training process, optimizes the architecture by removing unnecessary components, and broadens the framework’s applicability across multiple domains and modalities, including challenging OCR-intensive tasks. This adaptation results in a diverse suite of models that vary in parameter size and support multilingual capabilities.
LAVIS [253] was presented as an open-source deep learning library tailored for language-vision integration. LAVIS provides a unified platform facilitating access to cutting-edge image and video language models, along with a suite of tools for a wide range of multimodal tasks. The library is designed for scalability and flexibility, preparing it for future expansion and customization, with comprehensive benchmarking results demonstrating its effectiveness in language-vision applications.
V Outlook and Future Opportunities
V-A Data Generation
Generative models such as the Diffusion Model [91] have achieved significant success in NLP and computer vision. However, their application in industrial settings remains underexplored. In domains like equipment fault diagnosis and anomaly detection, where small sample sizes are a frequent challenge, the practical implementation and effectiveness of these models require further investigation. The integration of generative models into these specific industrial applications holds potential, but their suitability for handling limited data scenarios is still an open question.
As the demand for AIGC continues to grow, deploying AIGC within mobile edge networks presents a critical area for exploration [7]. These networks offer real-time, personalized AIGC services while ensuring user privacy. However, advancements in supporting technologies and the expansion of practical application scenarios are needed to fully realize the potential of AIGC in these decentralized environments. Addressing these challenges is essential to enable scalable and efficient deployment of AIGC in diverse industrial settings.
Finally, improving instruction adjustment techniques is crucial for enhancing large language models’ ability to interpret and respond accurately to instructions across broader scenarios. Current methods range from manual annotation to leveraging LLMs like the GPT series for automated generation of training data. Although initial results have been promising, integrating these methods into more efficient workflows remains a topic for future research. Furthermore, advancements in generative technologies such as Diffusion Models have increased the realism of synthetic data, including Deepfakes [254]. This creates significant challenges for verifying authenticity, both for human observers and machine classifiers. To mitigate the risks associated with Deepfakes, detection methods are being strengthened, incorporating spatial, frequency, and temporal analyses as well as techniques for identifying synthetic artifacts [255].
V-B Model/Knowledge Reuse
Model reuse employs strategies such as duplication [256], stacking [257] and combining [258] weights of pre-trained smaller models or using knowledge distillation to initialize larger models. The fundamental intuition is that reusing smaller models to warm-start the training of large transformers provides a robust initial point, thereby stabilizing optimization from the outset and conserving data samples needed to explore optimization trajectories. The scaling laws [259] of models can be through model reuse reproduced, even of extreme data conditions scarcity.
Some studies enhance future learning by establishing a knowledge pool [114]. When adapting to new tasks, relevant knowledge is first retrieved from a knowledge base according to the specific distribution of the new task. Subsequently, the pre-trained model is fine-tuned using knowledge transfer methods, such as joint reparameterization, to facilitate effective knowledge reuse. This approach enhances the model’s adaptability to new tasks and optimizes learning efficiency.
V-C Beyond Single Modality
Recent advancements are not confined to unimodal models but have expanded to multimodal pre-trained Models, such as CLIP [20]. These multimodal PTMs are capable of processing, responding to, and reasoning with various types of inputs. While significant progress has been made in visual recognition, particularly in using textual information to select appropriate prompts, interest in going beyond visual recognition is rapidly growing.
For example, to automate the generation of Electrocardiogram reports and ensure their versatility, a Multimodal Electrocardiogram IT framework has been proposed, attempting to utilize LLMs and multimodal instructions to generate Electrocardiogram reports [260]. ImageBind [261] presents a method for joint embedding across six different modalities—images, text, audio, depth, thermal, and IMU data. It extends its zero-shot capabilities to new modalities just by leveraging its natural pairing with images. It enables novel out-of-the-box applications including cross-modal retrieval, composing modalities with arithmetic, cross-modal detection, and generation. These studies mark a significant step forward in multimodal learning, with the integration of multiple modalities opening new possibilities and pathways for future industrial research and applications.
V-D Learning with Restricted Resources
LMs have demonstrated substantial efficacy in industrial applications, yet their continual tuning often demands significant computational resources. The deployment of these models extends beyond cloud-based environments to edge devices, including FPGAs, where models are often quantized and hard-coded to accelerate processing, thus rendering them non-modifiable. Recent advances in CL research, as indicated by Prabhu et al. [262], increasingly focus on applying models in resource-constrained scenarios. This trend is expected to highlight and address key challenges related to computational efficiency in future developments.
Traditional methods typically rely on computationally intensive backpropagation for model updates, which may not be feasible on some devices. As an innovative solution, the Forward-Only Adaptation [263] method adapts models solely at test time using a derivative-free covariance matrix evolution strategy, circumventing the need for backpropagation. This strategy provides a new possibility for model adaptation in environments with limited computational resources.
Edge-based LMs can achieve fast, near real-time analysis and response; however, due to the limited resources of edge nodes and end devices, they struggle to support such large models [34, 264]. Therefore, it is essential to integrate end, edge, and cloud computing to fully leverage their respective advantages, thereby compensating for the resource limitations at any single level.
VI Conclusion
This paper comprehensively surveys and forecasts the development prospects of IIoT-Enhanced General Industrial Intelligence based on large models. Initially, the paper outlines typical application scenarios in the industrial sector and introduces industrial-scale models with significant potential. The discussion is then divided into two main themes: the technological support provided by large models (LMs for GII) for general industrial intelligence and the potential challenges these models face in actual industrial applications (LMs on GII). We explore how pre-trained LMs can be developed to address specific industrial needs and discuss how CL and incremental learning can further enhance the adaptability and sustainability of these models.
Lastly, the paper discusses future research directions, including data generation, model reuse, and multimodal learning. This review aims to inspire more practitioners and researchers to explore continual learning to advance the development of robust and self-evolving industrial AI systems.
VII Acknowledgements
We would like to extend our special thanks to Ruyi Huang, Longya Xiao, Yihao Wang, Chenguang Ma, Xiang Zhang, Qiurui Zhang and Penghong Lu from the Shien-Ming Wu School of Intelligent Engineering, South China University of Technology, China, for their invaluable support in providing data and materials in the field of industrial applications during the writing of this paper.

### v2

@@ VI-C3 Parameter Isolation Methods
Parameter isolation methods address catastrophic forgetting by allocating specific parameters within a network for each task. These methods can be categorized into static architecture methods and dynamic architecture methods, depending on whether the network structure remains fixed during training.
Static Architecture Methods. Static architecture methods operate under a fixed network structure, allocating specific subsets of parameters for task-specific learning.
PathNet [223] introduces pathways that consist of active modules in each layer of the network. Each task is assigned its optimal pathway, which becomes fixed after training. This prevents further modifications to the parameters associated with that task, effectively preserving the acquired knowledge.
PackNet [224] leverages pruning techniques to allocate parameters for new tasks. Initially, the network learns the parameters for the first task, after which a subset of the learned weights is pruned (set to zero). The pruned weights are then reallocated and trained for subsequent tasks. For each task $t$, after training, pruning is applied to the weights allocated for task $t$, and these pruned weights are retrained for task $t+1$. To preserve knowledge from previous tasks, weights associated with earlier tasks remain frozen. Masks are maintained to indicate the active parameters relevant to each task, enabling the network to selectively activate specific parameters during task inference.
HAT [225] introduces a hard attention mechanism that employs task-specific masks at the unit level. These masks constrain parameter updates during new task learning, ensuring the retention of knowledge from prior tasks while enabling task-specific learning.
Dynamic Architecture Methods. Dynamic architecture methods allow the network to expand as new tasks are introduced, retaining the parameters associated with previously learned tasks.
Progressive Neural Networks (PNNs) [226] address catastrophic forgetting by creating a new network for each task. Knowledge transfer is facilitated through lateral connections, which enable the new network to utilize features learned by earlier networks. This approach ensures the preservation of previously acquired knowledge while supporting task-specific learning for new tasks.
Dynamic Expandable Networks (DEN) [227] utilize group-sparsity regularization to determine the necessity of adding new neurons at each layer for a given task. By expanding the network only when required, DEN minimizes unnecessary network retraining and maintains efficiency.
VI-D Emerging Continual with LPMs
As shown in Fig. 27, the field of continual learning with large pre-trained models (LPMs) has evolved significantly in recent years, driven by innovative approaches designed to address the challenges of adapting to new tasks while preserving prior knowledge. This section introduces three prominent paradigms—Delta-tuning CL, Modular CL, and Analytic CL—and highlights their unique mechanisms and contributions.
Figure 27: Development timeline of large pre-trained models based continual learning.
VI-D1 Delta-tuning CL
Delta-tuning continual learning (Delta-tuning CL) is an emerging paradigm that emphasizes the principles of minimal modification and high efficiency. This method freezes most parameters of LPMs and introduces a small number of trainable parameters to adapt to new tasks [164, 229, 230, 231]. This design contrasts sharply with traditional full-parameter fine-tuning methods. Delta-tuning CL preserves the generalization ability of LPMs to the greatest extent while significantly reducing computational and memory costs. This approach is particularly suitable for resource-constrained scenarios and environments with complex and dynamic tasks [161].
VI-D2 Modular CL
Modular continual learning (Modular CL) is a novel framework that uses modular structures and routing mechanisms to represent and compose knowledge dynamically [232, 233, 234]. This framework divides knowledge into independent modules. Each module serves as a distinct functional or structural unit. The routing mechanism selects or combines modules dynamically based on the characteristics of the input. This design creates an adaptive system tailored to the requirements of specific tasks.
VI-D3 Analytic CL
Figure 28: Analytic CL Based on Pre-trained Models vs. Traditional Replay-based CL.
As shown in Fig. 28, Analytic continual learning (Analytic CL) is a continual learning paradigm centered on analytic methods, utilizing explicit computation of model parameter updates to achieve efficient retention of historical knowledge. Unlike traditional optimization methods based on gradient descent, Analytic CL employs linear algebra and analytical solutions to update model parameters.
The initial representative method, Analytic Class-Incremental Learning (ACIL) [235, 236] focuses on analytic learning by recursively updating weight matrices and regularized feature autocorrelation matrices (RFAuM) [237]. ACIL achieves theoretical absolute memorization by reconstructing results identical to joint learning using only data from the current task without requiring historical data.
As analytic methods advance, Random Projections and Class-Prototype Strategy (RanPAC) [238] integrates LPMs and feature space optimization, expanding the practical applications of Analytic CL. RanPAC employs frozen random projection layers and nonlinear random projections to map features from LPMs into higher-dimensional spaces, improving feature linear separability. Additionally, the method uses class-prototype strategies and second-order feature statistics, such as covariance matrices, to represent class distributions.
VII Outlook and Future GII
The IIoT is not merely a data pipeline but serves as an intelligent infrastructure for the development of LMs. With its vast data resources, low-latency high-bandwidth networks, and edge computing capabilities, the IIoT provides essential support for the training, deployment, and continuous evolution of LMs. As illustrated in Fig. 2, the lifecycle of LMs within the IIoT consists of four key stages. The following sections outline future research directions and technical advancements anticipated in each of these stages.
VII-A Data Generation in IIoT
Data generation, particularly diffusion models [108] and other generative AI techniques, are playing an increasingly critical role in IIoT. These methods can produce high-quality synthetic data to address data scarcity, enhance model robustness, and ensure data privacy.
Lesson Learnt: Diffusion models demonstrate great potential in small-sample scenarios, but their high computational cost requires further optimization to suit resource-constrained edge devices [111]. For time-series data [239] and image/video data [240], generation techniques must account for temporal and multimodal characteristics while incorporating domain-specific prior knowledge for better optimization. Moreover, data generation can simulate diverse fault scenarios, providing richer datasets for model training and evaluation.
VII-B Model Training in IIoT
Federated learning and continual learning are set to become the dominant approaches for model training in IIoT. Federated learning effectively addresses challenges such as data silos and privacy concerns, while continual learning enables models to adapt to dynamically changing environments. Recent research in CL increasingly focuses on applying models in resource-constrained scenarios. This trend is expected to further highlight computational efficiency challenges and propose solutions in future developments [241].
Lesson Learnt: The application of federated learning in IIoT faces challenges related to communication efficiency, model consistency, and security. Asynchronous federated learning, model compression, and blockchain technologies have been shown to effectively mitigate these issues. For continual learning, addressing catastrophic forgetting and knowledge transfer remains critical. While traditional replay- and regularization-based methods have limitations, modular and analytic approaches exhibit greater potential.
Edge-based LMs can achieve fast, near real-time analysis and response. However, due to the resource constraints of edge nodes and end devices, they struggle to support such large models [242, 243]. Therefore, integrating end, edge, and cloud computing to leverage their respective advantages is essential to compensate for resource limitations at any single level.
VII-C Model Routing and Merging in IIoT
As the scale and complexity of LLMs continue to grow, model routing and merging technologies are becoming increasingly important. These techniques enable the deployment of different model components across various devices and dynamically combine models based on task requirements, achieving more efficient and flexible model deployment and execution. For instance, the “AI Flow” framework dynamically evaluates task demands and network conditions to allocate inference tasks optimally, improving resource utilization [175]. On the other hand, model merging integrates models trained on different devices, combining diverse data and knowledge to significantly enhance model performance and generalization. For example, the TAKFL [244] framework employs independent distillation, self-regularization, and task-level arithmetic knowledge integration to achieve efficient and secure model merging.
Lesson Learnt: Dynamic task allocation and model integration can improve system efficiency and robustness; however, they require careful consideration of device heterogeneity and resource constraints to ensure optimal performance. Furthermore, context-aware routing and performance-based routing are two commonly used strategies that can enhance task assignment by dynamically adapting to environmental conditions and resource availability. Additionally, model merging must address challenges such as model conflicts and knowledge integration to effectively combine knowledge from diverse models while maintaining system consistency and accuracy.
VII-D Continuous Evolution in IIoT
The continual evolution of LLMs in the IIoT represents a critical area of research. Models must continuously acquire new knowledge and skills to adapt to ever-changing environments and task requirements. In terms of model reuse, IIoT provides distributed computational capabilities, enabling techniques such as duplication [245], stacking [246], and weight combination [247] to be efficiently implemented in resource-constrained environments. Additionally, IIoT supports knowledge distillation to initialize large Transformer models, significantly reducing the need for data samples and stabilizing optimization processes. Even under extreme data scarcity, model reuse can effectively reproduce scaling laws [248], ensuring efficient and reliable learning.
Lesson Learnt: Existing continual learning methods primarily focus on addressing catastrophic forgetting in classification tasks but need to be extended to other task types, such as regression, prediction, and control. Furthermore, new continual learning approaches must be developed to enhance learning efficiency, generalization, and interpretability. Zero-waste continual learning emphasizes maximizing the utilization of existing resources and computational capacity, making it a pivotal direction for the future of continual learning. For instance, through the use of a knowledge pool [164], models can rapidly access task-relevant knowledge to adapt to new tasks. By employing fine-tuning techniques such as joint reparameterization, pre-trained models can efficiently transfer and integrate this knowledge, optimizing learning efficiency and adaptability.
VIII Conclusion
This paper explores the evolution of the Industrial Internet of Things (IIoT) from a traditional data pipeline to an intelligent infrastructure that supports the full lifecycle of industrial large models. First, it introduces various types of industrial large models and their typical applications, focusing on fault diagnosis, predictive maintenance, planning, and control. Next, the paper examines the rich data resources provided by IIoT, highlighting how techniques such as data freshness, data denoising, selection, and generation ensure data quality to support model training and fine-tuning. It then analyzes the infrastructural role of IIoT in model training, emphasizing how low-latency and high-bandwidth networks enable efficient model learning and real-time adaptation. Subsequently, the paper explores the advantages of deploying large models in IIoT environments, particularly how modular design and routing mechanisms enhance system flexibility and adaptability to meet diverse industrial needs. Finally, the paper examines the continual learning mechanisms of large models, focusing on how traditional and pre-trained model-based continual learning methods ensure long-term effectiveness and adaptability in dynamic industrial environments. Through this comprehensive discussion, the survey provides a thorough perspective and in-depth analysis of the application of large models in IIoT and their contribution to advancing General Industrial Intelligence (GII).
In the future, we will focus on providing a more detailed exploration of the application scenarios of industrial large models in key stages such as research and development, production and manufacturing, testing and validation, and operational management, highlighting their potential value and practical applications.
IX Acknowledgements
We would like to extend our special thanks to Ruyi Huang, Longya Xiao, Yihao Wang, Qiurui Zhang, Chenguang Ma, Xiang Zhang, Di Fang and Run He from the Shien-Ming Wu School of Intelligent Engineering, South China University of Technology, China, for their invaluable support in providing case studies, data, and materials related to industrial applications, which greatly enriched the depth and practical relevance of this paper.